# Random Utility with Aggregation

### v2

@@ B.1.1 Effect of composition distributions $\lambda$
(a) Heatmap of Biases across $\lambda_{\{x,a_{0}\}}$.
@@
(a) Heatmap of Biases across $\lambda_{\{x,a_{0}\}}$.
@@
B.1.2 Maximum/Minimum, and independent case
@@
B.2 Effect of preference structure on bias
In this section, we repeat the simulation in Subsection 5.2.3 with different parameter values. The setup is the same as in Figure 3 but with different values of $\lambda$.
The figures show the biases across utility values of $z$ and $w$. In general, the bias is largest when overlappingness is most likely (the top left and bottom right corners of each square) and smallest when overlappingness is less likely (the top right and bottom left corners of each square), consistent with Conjecture (B) and Figure 3.
B.3 Distance to the ARU Polytope
B.3.1 Effect of composition distribution $\lambda$ on distance
In this section, we repeat the simulation in Subsection 5.3.1 with different parameter values. The setup is the same as in Figure 5 but with different values of $\lambda$.
@@ B.3.1 Effect of composition distribution $\lambda$ on distance
(a) Heatmap of Distance across $\lambda_{\{x,a_{0}\}}$.
@@
(a) Heatmap of Distance across $\lambda_{\{x,a_{0}\}}$.
@@
B.3.2 Effect of preference structure on distance
Finally we repeat the simulation in Section 5.3.2. These four figures are the same as Figure 6 with different values of $\lambda$ — they show the distance from the ARUM polytope across across utility values of $z$ and $w$.

### v3

@@ E.2.1 Effect of composition distributions $\lambda$
a Heatmap of Biases across $\lambda_{\{x,a_{0}\}}$.
@@
a Heatmap of Biases across $\lambda_{\{x,a_{0}\}}$.
@@
E.2.2 Maximum/Minimum, and independent case
@@
E.3 Effect of preference structure on bias
In this section, we repeat the simulation in Subsection 5.2.2 with different parameter values. The setup is the same as in Figure 2b but with different values of $\lambda$.
The figures show the biases across utility values of $z$ and $w$. In general, the bias is largest when overlappingness is most likely (the top left and bottom right corners of each square) and smallest when overlappingness is less likely (the top right and bottom left corners of each square), consistent with Conjecture (B) and Figure 2b.
E.4 Distance to the ARU polytope
E.4.1 Effect of composition distribution $\lambda$ on distance
In this section, we repeat the simulation in Subsection 5.3.1 with different parameter values. The setup is the same as in Figure 5a but with different values of $\lambda$.
@@ E.4.1 Effect of composition distribution $\lambda$ on distance
a Heatmap of Distance across $\lambda_{\{x,a_{0}\}}$.
@@
a Heatmap of Distance across $\lambda_{\{x,a_{0}\}}$.
@@
E.4.2 Effect of preference structure on distance
Finally we repeat the simulation in Section 5.3.2. These four figures are the same as Figure 5b with different values of $\lambda$ — they show the distance from the ARUM polytope across across utility values of $z$ and $w$.